Snakes shed their skin to enable growth. If the time it takes for a snake to shed its akin is normally distributed with mean of 8 days and standard deviation of 2 days, what is the probability that a snake will shed its skin after 7 days?

1
Expert's answer
2021-11-04T16:46:48-0400
"P(X>7)=1-P(X\leq7)"

"1-P(Z\leq\dfrac{7-8}{2})=1-P(Z\leq-0.5)"

"\approx0.691462"

The probability that a snake will shed its skin after 7 days is 0.691462.